OpenStudy (anonymous):

What is the equation of the inverse of h(x) = 5x + 2?

OpenStudy (btaylor):

y = 5x + 2 switch x and y, then solve for y: x = 5y + 2 x-2 = 5y y = (x-2)/5

Parth (parthkohli):

First, write it as \(y = 5x + 2\) Now get the value of x in terms of y, and then replace y with x and x with h(x)

Parth (parthkohli):

\( \color{Black}{\Rightarrow \large {y - 2 \over 5} = x }\) Replace \( \color{Black}{\Rightarrow h^{-1}(x) =\Large {x -2 \over 5} }\)